OpenStudy (gabylovesyou):

@misty1212 How can I find the equation of a line in standard form if the line passes from (0,3) to (-3,0) I got x + y = -3 correct ?

OpenStudy (misty1212):

lets see did you find the slope?

OpenStudy (misty1212):

ok good, so now we start with \[y=x+3\] right ?

OpenStudy (misty1212):

damn

OpenStudy (misty1212):

when you see the point \((0,3)\) you should think right away "y intercept is 3"

OpenStudy (misty1212):

ok fine \[y-3=1(x-0)\] so \[y-3=x\] right

OpenStudy (misty1212):

how did that - 3 get to the other side of the equal sign? jump?

OpenStudy (misty1212):

ok but how is \(0+3=-3\)??`

OpenStudy (misty1212):

when you add 3, you add 3! it doesnt become minus

OpenStudy (misty1212):

there are no parentheses, you are not distributing anything, you are just adding 3

OpenStudy (misty1212):

\[7-3=4\\ 7=4+3\]

OpenStudy (misty1212):

\[y-3=x\\ y=x+3\]

OpenStudy (misty1212):

btw there is no such thing as \(-0\) as soon as you see zero, dont write it \[y-3=1(x-0)\] is the same as \[y-3=x\]

OpenStudy (misty1212):

now do you need to write it in standard form as well?

OpenStudy (gabylovesyou):

yes

OpenStudy (misty1212):

yes, point slope is \(y-3=x\) slope intercept is \(y=x+3\)

OpenStudy (misty1212):

then if you want to write it in standard from you want all the variables on one side, the number on the other you could subtract x and write \[-x+y=3\] or if want to make the coefficient of the x term positive, write \[x-y=-3\]

OpenStudy (gabylovesyou):

i thought u couldnt subtract x ?

OpenStudy (misty1212):

you can add, subtract whatever you just can't "bring it across" the equal sign

OpenStudy (misty1212):

\[y=x+3\] subtract \(x\) from both sides you get \(y-x=3\)

OpenStudy (misty1212):

then multiply both sides by \(-1\) and get \(x-y=-3\)

OpenStudy (gabylovesyou):

what do u mean by "bring it across? and i prefer -x + y = 3

OpenStudy (misty1212):

nvm about "bring it across" if you prefer \(-x+y=3\) then leave it as that one !

OpenStudy (gabylovesyou):

lol okay ;) Thank you (again)

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]